OpenStudy (anonymous):

In 1991, the cost of mailing a 1 oz. first-class letter was 29 cents, and the inflation rate was 4.6%. If the inflation rate stayed constant, the function C(t) = .29(1.046)t would represent the cost of mailing a first-class letter as a function of years since 1991. Question: In 2007, the cost of mailing a first-class letter was 41 cents. Has the inflation rate stayed constant since 1991? Explain.

OpenStudy (mathstudent55):

You mean C(t) = 0.29(1.046)^t, right? t is an exponent, isn't it? If so, then t = years since 1991, so 2007 - 1991 = 16 C(16) = 0.29(1.046)^16 = 0.60 Answer is no.

OpenStudy (anonymous):

yes it is an exponent :) thank you so so much!! @mathstudent55

OpenStudy (anonymous):

do you know this question having to do with the same problem? 1. If the function given holds true, in what year would the cost of mailing a first-class letter reach 60 cents? @mathstudent55

OpenStudy (anonymous):

:)

OpenStudy (mathstudent55):

When we let t = 16, which represents 1991 + 16 = the year 2007, we did get C = 0.60, so the answer is 2007.